OpenStudy (anonymous):

each interior angle of a regular polygon is four times the measure of each exterior angle. How many diagonal does the polygon have?

OpenStudy (anonymous):

Same game as before....

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

i need a complete answer. plss help me..

OpenStudy (anonymous):

((2x-4)*90)/x=4(360/x) solve it u will get the number of sides the decide the number of diagonal

OpenStudy (anonymous):

x + 4x = 180 -> x = 36 so 10 sides.

OpenStudy (anonymous):

Number of diagonals is n/2(n-3) where n is number of sides.

OpenStudy (anonymous):

n/2*(n-3)

OpenStudy (anonymous):

the answer is 56?

OpenStudy (anonymous):

How u get that?

OpenStudy (anonymous):

i mess its 35

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

ur welcome.